Jamison Kovach Supply Company manufactures paper clips and other office products. Although​ inexpensive, paper clips have provided the firm with a high margin of profitability. Sample size is 40. Results are given for the last 10 samples Sample 1 2 3 4 5 6 7 8 9 10 Defectives 4 7 3 6 6 5 4 4 2 12 Establish the control limits to include​ 99.73% of the random variation in defectives. UCL= LCL= Has the process been in​ control? Based on the developed control​ limits, the number of defectives has been?
Answer #1

Answer: 99.73% explains the requirement of 3 sigma control limits

Day Defective Observations proportion of defect= defectives/observations
1 4 40 0.100
2 7 40 0.175
3 3 40 0.075
4 6 40 0.150
5 6 40 0.150
6 5 40 0.125
7 4 40 0.100
8 4 40 0.100
9 2 40 0.050
10 12 40 0.300
total 53 400
steps and formulas
Probability of defects=P= total defectives /total observations 0.1325
Q= 1-P 0.8675
N= average sample size 40
UCL= P + 3*squareroot(P*Q/N) 0.293
LCL= P - 3*squareroot(P*Q/N) -0.0283
defects cannot be negative, therefore negative LCL is taken as '0' 0.000
Standard deviation squareroot(P*Q/N) 0.0536
  • 10th sample is out of control. As the proportion of defectives is 0.300 in 10th sample and UCL is 0.293. so, 10th sample is beyond UCL
  • Total no. of defectives have been 53.

Do rate the answer